ID |
Date |
Author |
Type |
Category |
Subject |
943
|
Wed Aug 18 03:01:23 2010 |
rana | Laser | GYRO | Excess noise mystery resolved |
After some navel gazing, this all no longer seems so surprising. A block diagram of the system with the noise terms labeled should reveal why there is almost no beat signal with the CW servo turned off. I believe that there is no gyro signal in the modulation off case.
|
942
|
Tue Aug 17 15:34:02 2010 |
Alastair | Laser | GYRO | Changed to Marconi on AOM |
Quote: |
Its amazing if it turns out that the noise goes down that much with just turning off the feedback on the CW path. Is it really repeatable, or is it just that the CW beam was not resonating and therefore not producing a beat signal in the transmisssion?
|
I have repeated the measurement here. I began by measuring the transmission signal on an analyzer with the modulation on and off. With it on we're getting -26.5dBV at 95.021MHz. With the Tektronix set to continuous we're getting -26.6dBV at 95.019MHz.
I then repeated the noise measurement with the PLL to check that I didn't make a mistake yesterday. I made sure both directions looked locked on the CRT while taking the data and afterwards plugged the output back into the analyzer and measured the peak to be -27.1dBV. The plot is below. We're still seeing the large drop in noise with the modulation off. |
Attachment 1: gyro_transmission_08_17_10.pdf
|
|
941
|
Tue Aug 17 10:18:02 2010 |
rana | Laser | GYRO | Changed to Marconi on AOM |
Its amazing if it turns out that the noise goes down that much with just turning off the feedback on the CW path. Is it really repeatable, or is it just that the CW beam was not resonating and therefore not producing a beat signal in the transmisssion?
We need to look at the noise of the PD, the PDH box, and the phase noise of the AOM drivers and then put them into the loop model and plot them with the noise budget. Do we have data on the open loop gain for the CCW and CW?
I also suggest that instead of using DTT, you make a mDV/NB hybrid that can plot each day's noise curve onto our latest SVN NB model. It would be handy. |
940
|
Mon Aug 16 19:01:20 2010 |
Alastair | Laser | GYRO | Noisy noise stuff |
I've done a couple of things since the earlier post, none of which have reduced the noise level. First I put some lead bricks (and a couple of SR560s once I ran out of lead bricks) on top of the cover. I put a sheet of foam between these and the lid, in the vain hope that this might damp some accoustic coupling through the top.
Next I put in two beam dumps behind the flat mirrors in the cavity. These should really have been in already since there is a small amount of transmission through them. I can't get to the other two transmitted beams because of the mirror mounts. We might need mounts that allow access to all beams in and out in order to block these.
I went back to looking at the locking signals again. The slow feedback on the CCW loop appeared to be oscillating a bit and I turned the gain down till it stopped - a factor of 10 below where I started. I also added the boost on this side.
I then cranked up the gain on the CW loop as well, and added the boost. I turned it up till it started oscillating and then came back down by a small amount. Still no different.
I find it intriging that the noise is much lower with the FM modulation turned off to the AOM. It seems to me that this might point to it coming either from the locking loop for the AOM, or a mechanical noise source, since it doesn't appear to be the Tektronix at this level. |
939
|
Mon Aug 16 18:56:28 2010 |
1954 B-movie | Lab Infrastructure | ANTS! | How do we get rid of ants? |

Quote: |
Quote: |
It's that time of year again in Pasadena and the ants are going a bit crazy. There are about 10 on the floor in the ATF, and no doubt there will soon be about 1e6 of them. Do we have a standard ant removal scheme for cleanrooms? The stuff I use at home is a spray and probably not very nice for optics. We could use some of the slightly illegal ant chalk I guess.
|
Steve has the cleanroom janitor mop the 40m floor in the LVEA/IFO room with diluted ant poison.
|
|
938
|
Mon Aug 16 18:36:59 2010 |
Jenne | Lab Infrastructure | ANTS! | How do we get rid of ants? |
Quote: |
It's that time of year again in Pasadena and the ants are going a bit crazy. There are about 10 on the floor in the ATF, and no doubt there will soon be about 1e6 of them. Do we have a standard ant removal scheme for cleanrooms? The stuff I use at home is a spray and probably not very nice for optics. We could use some of the slightly illegal ant chalk I guess.
|
Steve has the cleanroom janitor mop the 40m floor in the LVEA/IFO room with diluted ant poison. |
937
|
Mon Aug 16 15:31:27 2010 |
Zach | Laser | GYRO | Changed to Marconi on AOM |
It is so great that we have a working readout now so we can monitor our sensitivity in real time. Now that we have this, we have to systematically go through and make each component quieter until we get sensitive enough to see one of our expected sources. When I get back I'll hook up with Rich to get started on the RFPDs. I am betting that a huge amount of our noise is coming from their absence. We should also have a think about if and how we need to modify the PDH boxes. I am a little out of the loop by now, but it will be nice to get back into the swing of things.
I gave a lunch talk about the gyro here at LHO today, and I was happy to report our getting within 106 of design sensitivity ;-)
Quote: |
I borrowed a second Marconi from Frank this morning to check whether the Tektronix driving the AOM was our main noise source. As you can see from the plot the noise was virtually identical using the Tektronix. I also (accidentally) took data with the modulation input on the Marconi switched off, but the carrier on (so the CW loop was resonant in the cavity still) - interestingly the noise was quite a lot lower - I need to think a bit more about this, but it gives an indication that certain noise sources are not the main ones at the moment (certainly the Mach - Zehnder readout isn't the dominant noise source). I also took the lid off the box and remeasured - the noise below 3Hz was worse. It is up about a factor of 10 at 0.01Hz. Nice to know our box is doing something for us.
I'm going to pass this Marconi back to Frank now. I'll try to find a replacement from the 40m if there is a spare one after the phase measurements.
|
|
936
|
Mon Aug 16 14:19:00 2010 |
Alastair | Laser | GYRO | Changed to Marconi on AOM |
I borrowed a second Marconi from Frank this morning to check whether the Tektronix driving the AOM was our main noise source. As you can see from the plot the noise was virtually identical using the Tektronix. I also (accidentally) took data with the modulation input on the Marconi switched off, but the carrier on (so the CW loop was resonant in the cavity still) - interestingly the noise was quite a lot lower - I need to think a bit more about this, but it gives an indication that certain noise sources are not the main ones at the moment (certainly the Mach - Zehnder readout isn't the dominant noise source). I also took the lid off the box and remeasured - the noise below 3Hz was worse. It is up about a factor of 10 at 0.01Hz. Nice to know our box is doing something for us.
I'm going to pass this Marconi back to Frank now. I'll try to find a replacement from the 40m if there is a spare one after the phase measurements. |
Attachment 1: gyro_transmission_08_16_10.pdf
|
|
935
|
Mon Aug 16 13:25:03 2010 |
Zach | Lab Infrastructure | ANTS! | How do we get rid of ants? |
Lighter fluid and a torch; the ol' "line of fire" method always worked for me.
Quote: |
It's that time of year again in Pasadena and the ants are going a bit crazy. There are about 10 on the floor in the ATF, and no doubt there will soon be about 1e6 of them. Do we have a standard ant removal scheme for cleanrooms? The stuff I use at home is a spray and probably not very nice for optics. We could use some of the slightly illegal ant chalk I guess.
|
|
934
|
Mon Aug 16 12:33:08 2010 |
Alastair | Lab Infrastructure | ANTS! | How do we get rid of ants? |
It's that time of year again in Pasadena and the ants are going a bit crazy. There are about 10 on the floor in the ATF, and no doubt there will soon be about 1e6 of them. Do we have a standard ant removal scheme for cleanrooms? The stuff I use at home is a spray and probably not very nice for optics. We could use some of the slightly illegal ant chalk I guess. |
933
|
Fri Aug 13 18:22:04 2010 |
Jenna | Laser | GYRO | Successful PLL |
After we got both sides locking properly and not oscillating anymore (as I put in the last post), we swapped from using the Marconi to drive the AOM back to the Tektronix (because we only have one Marconi), and immediately the noise in the spectrum analyzer looked worse. We took this signal and mixed it with the Marconi, and with the modulation switched off on the Tektronix, we looked at the beat frequency on a scope and minimized it. Then we turned on the PLL feedback to the Marconi, and now the error signal is a flat line. We turned the modulation on again to the Tektronix, and we can see a varying output signal which is our gyro signal, which is plotted below.
The calibration for this plot is 6.103e-4 V/count * 100kHz/V * lambda*S/(4A) * 2pi (rad/s/Hz) = 5.1686e-4.
DYM - Handshake: If you open the plot in preview (on a mac), and then save it as a pdf over itself, it associates a png file with it automatically, which acts as a thumbnail. |
Attachment 1: gryo_transmission.pdf
|
|
932
|
Fri Aug 13 15:50:39 2010 |
Jenna | Laser | GYRO | Beam overlap etc |
Quote: |
The error signal from the CW side looked much improved after changing the polarization (TEK00001.PNG) but the CW mode still looked unstable. When we lock both directions of the cavity and look at these signals again we see this (TEK00002.PNG) oscillation in the signal. It is massive. It basically takes up the full error signal and is no doubt the reason why the mode looks unstable.
If you turn off the FM input on the VCO driving the AOM then the oscillation is 50% smalller but still there. We wondered if this was coming from the Tektronix, so we swapped the AOM over to the Marconi, but we see exactly the same noise with this if the FM deviation is set to ~300kHz. It does go lower if we lower the deviation setting on the Marconi below about 100kHz. This might be the noise we're seeing in the 95MHz signal in transmission.
**Edit** we just spoke to Frank who thinks this may be because we have the gain too high in this loop (though turning down the PDH box to almost zero didn't get rid of the oscillation). Tomorrow we'll try reducing this further.
|
We were investigating the source of the oscillation in the CW signal when we found that the CCW photodiode was saturating, which must have happened when we were adjusting the power between the two beams. We put in a beamsplitter to reduce the power going onto the PD, and now the CW error signal is no longer oscillating so that seems to have been the source of the problem.
Our 95MHz signal is now at -28dBV, which is 1.3uW of laser power out of 42.5uW (3% contrast ratio), and the DC signal from the photodiode is 30mV which corresponds to 40uW of power on the PD out of 42.5uW, so we're getting most of the power onto the PD now. |
931
|
Fri Aug 13 10:18:20 2010 |
Jenna | Laser | GYRO | Pictures |
It's fixed now, in the original entry. |
930
|
Fri Aug 13 00:11:46 2010 |
Jenna | Laser | GYRO | Pictures |
Sorry, the input/output mirrors are flat, and it's the other two mirrors that are curved with a ROC of 9m. I'll try to fix it on the picture tomorrow. |
929
|
Thu Aug 12 23:26:06 2010 |
Zach | Laser | GYRO | Pictures |
Nice shots. I noticed that on one picture you indicate a "curved input mirror". Was the input mirror really changed to a curved one? If so, why? My (perhaps-flawed) intuition tells me that it will be more difficult to get an exact modematch to the weird phasefront on that mirror.
|
928
|
Thu Aug 12 22:55:22 2010 |
Jenna | Laser | GYRO | Pictures |
I've taken some pictures and tried to label the beam paths. They're not great, but hopefully they're understandable.
Green is the unsplit laser beam, Red is the CCW beam, Blue is CW (on the AOM side) beam, and Purple is both beams together.
The first three show the incoming and reflected beams. Arrows indicate a reflected beam. The last picture shows the beams in transmission onto the PD.
*Edited to correct first picture |
Attachment 1: Beam_Path.png
|
|
Attachment 2: beam_path_3.png
|
|
Attachment 3: beam_path_1.png
|
|
Attachment 4: beam_path_2.png
|
|
927
|
Thu Aug 12 18:32:37 2010 |
Alastair | Laser | GYRO | Beam overlap etc |
We started out today by trying to improve the beam overlap. After talking to Aidan, and with Koji and Rana's suggestions, we began by beating the beams on the BS at the output. We replaced the lens before the PD with a ~3cm focal length one. Then we used the adjuster on the BS itself to make the beams overlap at the PD. We got the signal to increase up to -31dBV (this was about 2% contrast).
We tried to improve this by adjusting the overlap by altering the BS, and one of the steering mirrors before the BS, but we couldn't improve it much.
Based on the 20mV signal from the DC input it seemed like we were only getting 27uW out of our 50uW on the actual photodiode. It is difficult to get the lens exactly the correct distance from the PD because we have to focus down so small (it's impossible to see how big the beam is using an IR card or viewer). We tried mounting the lens on a translation stage and maximising the power on the PD. We managed to get about 60% increase in power doing this.
Looking at the 95MHz signal on the spectrum analyzer the frequency shifts back and forward quite a bit (~100kHz). This goes away if you turn off the frequency modulation input on the VCO driving the AOM. Looking at the CW beam (the AOM direction beam) on the monitor the mode appears to be a little unstable even though the CCW beam is not. At this point we started looking at the reflection locking path for the AOM. The PD for this locking scheme was getting a bit saturated so we installed an 80% BS to reduce the power on the PD down to about 1.5mW.
At this point we looked at the errror signal from the CW direction while sweeping the laser. It looked a bit strange (see TEK00000.PNG below. The blue trace is the PD signal and the yellow is the error signal) and after making sure the beam was on the PD correctly we decided to check the polarization of the beam going into the cavity. We found that it had become slightly rotated. It probably got altered when we were aligning the beam through the Faraday. We rotated it back to S and checked that the cavity was still locking. We put a power meter on the transmission side and maximized the power transmitted by steering the beam into the cavity. The max we got was CCW 2.21mW out from 7.4mW in. We checked the CW polarization which was still correct, and also maximized it's power out (this side does not have a good mode matching solution in place) and we are getting 4.5mW our of 25.9mW. Since the output beams were now a little uneven in power we rotated the polarizer that controls the relatvie power in the two arms such that we are now getting ~12uW of power on the PD from each direction.
The error signal from the CW side looked much improved after changing the polarization (TEK00001.PNG) but the CW mode still looked unstable. When we lock both directions of the cavity and look at these signals again we see this (TEK00002.PNG) oscillation in the signal. It is massive. It basically takes up the full error signal and is no doubt the reason why the mode looks unstable.
If you turn off the FM input on the VCO driving the AOM then the oscillation is 50% smalller but still there. We wondered if this was coming from the Tektronix, so we swapped the AOM over to the Marconi, but we see exactly the same noise with this if the FM deviation is set to ~300kHz. It does go lower if we lower the deviation setting on the Marconi below about 100kHz. This might be the noise we're seeing in the 95MHz signal in transmission.
**Edit** we just spoke to Frank who thinks this may be because we have the gain too high in this loop (though turning down the PDH box to almost zero didn't get rid of the oscillation). Tomorrow we'll try reducing this further. |
Attachment 1: TEK00000.PNG
|
|
Attachment 2: TEK00001.PNG
|
|
Attachment 3: TEK00002.PNG
|
|
926
|
Wed Aug 11 22:42:52 2010 |
Koji | Laser | GYRO | attempted PLL |
I didn't get the meaning of "both beams are on the PD".
Yes, it is a required condition to get the beating. But, it is not the all.
As the beating is an interference of two beams, what you need is not "both beams on the PD" but "good overlap of two beams" (then put this combined beam to the PD).
i.e. You need to match the spots on the BS as well as on the PD.
Quote: |
Here is a summary of what we tried just now. To cut to the end of the story, we finally found that the 95MHz signal has disappeared from our PD. We don't know why because both beams are still on the photodiode.
Okay, back to the begining. We started by looking at the 95MHz signal on a spectrum analyzer and trying to maximize it by moving the beams on the PD. We weren't able to get the signal above ~ -68dBV. We increased the laser power on the PD to around 50uW, and we had a 95MHz signal of around -62dBV.
At this point we connected the signal to a mixer, and used a Marconi as the LO for the mixer. We set the carrier freq to match the AOM freq, and then set the Marconi to Ext FM, with deviation of 100kHz. We connected the output of the mixer to an SR560 which we connected back to the FM input of the Marconi. We were looking at the output of the mixer and the output of the SR560 on a scope. After quite some time we hadn't seen anything that made a lot of sense. We re-measured the 95MHz signal on the spectrum analyzer, and it had gone. We checked that both beams were on the PD, and that blocking either caused the voltage out to drop.
|
|
925
|
Wed Aug 11 18:51:13 2010 |
Alastair | Laser | GYRO | attempted PLL |
Here is a summary of what we tried just now. To cut to the end of the story, we finally found that the 95MHz signal has disappeared from our PD. We don't know why because both beams are still on the photodiode.
Okay, back to the begining. We started by looking at the 95MHz signal on a spectrum analyzer and trying to maximize it by moving the beams on the PD. We weren't able to get the signal above ~ -68dBV. We increased the laser power on the PD to around 50uW, and we had a 95MHz signal of around -62dBV.
At this point we connected the signal to a mixer, and used a Marconi as the LO for the mixer. We set the carrier freq to match the AOM freq, and then set the Marconi to Ext FM, with deviation of 100kHz. We connected the output of the mixer to an SR560 which we connected back to the FM input of the Marconi. We were looking at the output of the mixer and the output of the SR560 on a scope. After quite some time we hadn't seen anything that made a lot of sense. We re-measured the 95MHz signal on the spectrum analyzer, and it had gone. We checked that both beams were on the PD, and that blocking either caused the voltage out to drop. |
924
|
Wed Aug 11 14:36:54 2010 |
Alastair | Laser | GYRO | PD in transmission |
Quote: |
Let's get some overhead photos of the table with lots of the parts and beampaths labeled please. It will help in understanding these results so far.
For example, based on the peak height in the spectrum analyzer and the transmission diode's transimpedance gain, what is the contrast ratio for the two beams?
And let's get that signal demodulated with a Rubidium Marconi and fed into the DAQ! We should be able to get some long term data with it.
|
The peak on the spectrum analyser is at -68dBV. Converting this to volts (10^(-68/20)) we get 0.00040V
The transimpedance gain for the New Focus 1811 is 40V/mA and the response is approximately 0.75A/W. This means we have about 10^-9 amps of photocurrent, and 0.013uW of laser power.
The laser power on the photodiode is 16uW, so we have a contrast of about 0.08%......which doesn't seem like a lot. Admittedly our beams are different sizes coming out of the cavity and they are going through just one lens to focus down onto the PD, so perhaps this is where most of the contrast is lost. The beams are approximately the same power. We can increase the power on the PD a little as well. |
923
|
Tue Aug 10 23:44:17 2010 |
rana | Laser | GYRO | PD in transmission |
Let's get some overhead photos of the table with lots of the parts and beampaths labeled please. It will help in understanding these results so far.
For example, based on the peak height in the spectrum analyzer and the transmission diode's transimpedance gain, what is the contrast ratio for the two beams?
And let's get that signal demodulated with a Rubidium Marconi and fed into the DAQ! We should be able to get some long term data with it. |
922
|
Tue Aug 10 22:02:04 2010 |
Zach | Laser | GYRO | PD in transmission |

Quote: |
We put in the PD in transmission, and here is our lovely signal. We took the beam in transmission and put it through a beamsplitter and then a lens and onto the photodiode and aligned it by moving each beam one at a time and maximizing the DC signal from the PD. We then plugged the AC signal into a spectrum analyzer, and we can see the beat signal at 95MHz.
|
|
921
|
Tue Aug 10 20:29:53 2010 |
Jenna | Laser | GYRO | PD in transmission |
We put in the PD in transmission, and here is our lovely signal. We took the beam in transmission and put it through a beamsplitter and then a lens and onto the photodiode and aligned it by moving each beam one at a time and maximizing the DC signal from the PD. We then plugged the AC signal into a spectrum analyzer, and we can see the beat signal at 95MHz. |
Attachment 1: IGEJNS.TIF
|
|
920
|
Tue Aug 10 19:22:19 2010 |
Jenna | Laser | GYRO | Cavity is locking again! |
Upon Rana's suggestion, Alastair adjusted the polarization of the beam going through the EOM using the half wave plate before the EOM and a polarizing beamsplitter to make sure it was exactly S-polarized and locked the half wave plate at the correct position. He also tightened up some of the SMA cables, some of which were a bit loose. As you can see there's no more DC-offset in the PDH error signal (blue). |
Attachment 1: TEK00006.PNG
|
|
919
|
Tue Aug 10 18:48:09 2010 |
Jenna | Laser | GYRO | Error signal investigations |
We've put a beam dump in the cavity to block the beam, and the yellow plot is the reflected PD. The first plot shows the scope DC-coupled, and the second shows it AC-coupled with the sensitivity turned up on that channel. We're getting about 20mV of RF in that channel out of a 500mV signal. |
Attachment 1: TEK00004.PNG
|
|
Attachment 2: TEK00005.PNG
|
|
918
|
Tue Aug 10 18:23:46 2010 |
Jenna | Laser | GYRO | Error signal investigations |
The cavity still isn't locking properly, so we're sweeping the cavity to have a look at the error signal. The first picture is with everything setup the same way as when we walked in today, and for the second we adjusted the length of a BNC cable and messed with the phase knob on the PDH box to get rid of the DC offset on the error signal.
Blue is the PDH error signal, Yellow is the reflection PD signal, and pink is the waveform we're using to sweep the cavity. |
Attachment 1: TEK00001.PNG
|
|
Attachment 2: TEK00003.PNG
|
|
917
|
Tue Aug 10 13:44:49 2010 |
Zach | Laser | GYRO | Zach is feeeling excluded from this lack of locking |
I've been using LIGO overhead to investigate my new beard invisibility cloak. It's working.
|
916
|
Mon Aug 9 20:48:38 2010 |
Dmass | Laser | GYRO | Zach is feeeling excluded from this lack of locking |
I see no beard. |
915
|
Mon Aug 9 18:36:14 2010 |
rana | Laser | GYRO | Zach is feeeling excluded from this lack of locking |

|
914
|
Mon Aug 9 00:46:48 2010 |
Frank | Computing | General | netgpib |
i copied the network gpib scripts from 40m to /caltech/scripts/netgpibdata/ |
913
|
Sat Aug 7 19:04:07 2010 |
Dmass | Laser | Doubling | Mode Matching for Ovens |
Here is a semi complete plan of my "what next" when I get back:
- Finish Arm Mode Matching
- Relock PMC / spend a couple minutes fidding with gain and looking at ERR on a scope
- Tweak mirrors to reoptimize power output on a power meter
- Quick check of power in PMC input / refl / trans to see if loss is insane
- Block each arm and add a lens after the oven to make the spot (~500 um diam?) large at stolen ISS PDs
- Check that the size is still fine @ the green PDs
- Put Green PDs into my electronics box (that's where their transimpedance amps are)
- Align and check contrast defect. Pic of PDs on scope with PZT sweeping - verify
- Get more Green PD batteries just in case (maybe before 1 if I have to order)
- Rotate the ovens so the crystal axis is as close as I can make it to the optic axis. Find the phase matching temperature maximum crudely
- Move the ovens around along the axis to try to get a maximum power
- Get a good measurement of the phase matching curve: Block each arm and get the phase matching curve for each oven (~60 second time constant, so ~1-2 hour meas to get it well). Note this is only good for one angle, so every time I touch (re: twist) the oven, this phase matching width / temperature can change - I should be able to estimate this effect with less than a couple hours thinking). Note that this will be my theta/T calibration, and really needs to be a 5-10 percent number or so.
- Balance PDs (May have to replace op amps if any are bad at this point)
- Sweep the PZT of the MZ
- Look at Lissajous figure of Diff 1 vs Diff 2 with all 4 PD inputs connected, balance this with the pots inside the PD box
- Look at the Lissajous, make its crossing zero (so we can have Green Diff and IR Diff both be ~ zero)
- Get Min / Max of Lissajous for calibration
- Attempt to lock the Mach Zehnder (by first turning down the MZ gain, then increasing once lock is acquired)
- Post calibrated phase noise spectum to the elog
- ASSESS AT THIS POINT WHETHER TO DO STEPS 9 AND 10 OR JUST 9...OR EVEN JUST USING THE POOR PRECISION THERMOMETERS TO ESTIMATE PHASE NOISE FROM TEMPERATURE
- Calibrate the PZT Drive signal?
- Make a breakout cable from the front of the HV PZT Drive Box (On the PZT VMON output - D9 to BNC)
- Sweep PZT, look at ERR (phase noise difference, already calibrated)
- Have calibrated control signal now!
- Add thermometer to ADC
- Use the silver epoxy to directly glue temperature sensors from Frank to the oven
- Power them with an AD587 in series with some resistor (not really a current source, but it's fine)
- Calibrate thermometer using temperature sensor already on oven (RTDs), readout by the Newport 3040 temperature controller
- Add to ADC as calibrated readout of each oven temperature
- Subtract!?!?
- Look at coherence between temperature and phase noise in green (I expect there to be some)
- Do frequency domain subtraction of Temperature from Green Phase Noise with the scripts which I already wrote (this is easy)
- Compare the suppressed level of Green phase noise with the servod IR phase noise
- Compare suppression level to known noise floor of experiment. Figure out how best to present this plot
- ????????
- Profit.
|
912
|
Sat Aug 7 13:45:52 2010 |
Dmass | Lab Infrastructure | General | WHERE IS THE WINCAMD? |
Quote: |
Quote: |
I was trying to understand why the mode matching through the PMC was so bad (I got 30 mW transmitted with 120 mW input).
- I looked at the mode on the PMC REFL camera, and it did not look very guassian.
- I changed the laser current from 2 A to 2.2 A (what Rick said they ran it at at LHO) and saw no change in the mode
- I tried to take a beamscan...but could not find the WINCAMD and its computer (again.) I checked the usual suspects (PSL/TCS) and didn't see it.
So. Who knows where the wincamD is?
added: pics of PMC REFL before and after locking, in order
Totally possible that my mode matching is just that crappy? Maybe.
I did take beam scans of the mode going into the PMC just before aligning to it for the final time. I can do so again to get real profiles and characterize the H1NPRO
|
Here are some scans of the beam with the winCAMD. I had a reflective ND filter on for these (hence some of the fringing).
- 5-10 inches in front of NPRO (through mirror transmission)
- 21 inches in front of NPRO (maybe junk b/c of ND filter? Maybe not?)
- Downstream of: HWP and PBS to get 120 mW total (about 8% power), and some focusing lenses (this was at PMC waist). I think this is the "junk light" you get from an NPRO. Maybe this is a lot of power to have in junk, maybe not?
|
I matched this spatial mode to the cavity, and checked that it had the right waist size / location (5-10% on waist size, 1-2 inches on position) with the winCAMD. I was only able to get 220 mW out of the PMC after aligning tilt / translation into the PMC, with the full 1.6ish W in. At Rana's gentle suggestion, I am proceeding with the rest of the experiment. I noticed the PMC mirrors has a lot of scatter while I was aligning them, so I should quickly check if this is loss (Input = Refl + Trans + Loss?) by measuring Refl and Trans
Remember: If someone (myself?) later figures out what the problem is, I don't have to realign anything downstream of the PMC, I simply have to make there be more power coming out. |
911
|
Sat Aug 7 13:45:13 2010 |
Frank | Lab Infrastructure | General | wireless GPIB network devices now available |
i finished configuring the wireless bridges and the gpib-to -network adapters. The adapter have ip-addresses from 10.0.1.40 to 43.
All of them are online right now, but only the device with IP 10.0.1.43 is physically connected to an intrument right now.
The SR785 in the PSL lab has the IP 10.0.1.43, the other ones have to be installed at the instrument.
Once we've done that we should come up with some names for those devices, e.g. SR785-PSL, SR785-ATF or so, so that we do not have to remember the ip addresses all the time.
I hope we can use the names with the script too, i didn't try it so far. |
910
|
Fri Aug 6 16:56:58 2010 |
Aidan | Lab Infrastructure | General | WHERE IS THE WINCAMD? |
Looks like Hurricane Dmass.
Quote: |
Quote: |
I was trying to understand why the mode matching through the PMC was so bad (I got 30 mW transmitted with 120 mW input).
- I looked at the mode on the PMC REFL camera, and it did not look very guassian.
- I changed the laser current from 2 A to 2.2 A (what Rick said they ran it at at LHO) and saw no change in the mode
- I tried to take a beamscan...but could not find the WINCAMD and its computer (again.) I checked the usual suspects (PSL/TCS) and didn't see it.
So. Who knows where the wincamD is?
added: pics of PMC REFL before and after locking, in order
Totally possible that my mode matching is just that crappy? Maybe.
I did take beam scans of the mode going into the PMC just before aligning to it for the final time. I can do so again to get real profiles and characterize the H1NPRO
|
Here are some scans of the beam with the winCAMD. I had a reflective ND filter on for these (hence some of the fringing).
- 5-10 inches in front of NPRO (through mirror transmission)
- 21 inches in front of NPRO (maybe junk b/c of ND filter? Maybe not?)
- Downstream of: HWP and PBS to get 120 mW total (about 8% power), and some focusing lenses (this was at PMC waist). I think this is the "junk light" you get from an NPRO. Maybe this is a lot of power to have in junk, maybe not?
|
|
909
|
Fri Aug 6 14:12:22 2010 |
Dmass | Lab Infrastructure | General | WHERE IS THE WINCAMD? |
Quote: |
I was trying to understand why the mode matching through the PMC was so bad (I got 30 mW transmitted with 120 mW input).
- I looked at the mode on the PMC REFL camera, and it did not look very guassian.
- I changed the laser current from 2 A to 2.2 A (what Rick said they ran it at at LHO) and saw no change in the mode
- I tried to take a beamscan...but could not find the WINCAMD and its computer (again.) I checked the usual suspects (PSL/TCS) and didn't see it.
So. Who knows where the wincamD is?
added: pics of PMC REFL before and after locking, in order
Totally possible that my mode matching is just that crappy? Maybe.
I did take beam scans of the mode going into the PMC just before aligning to it for the final time. I can do so again to get real profiles and characterize the H1NPRO
|
Here are some scans of the beam with the winCAMD. I had a reflective ND filter on for these (hence some of the fringing).
- 5-10 inches in front of NPRO (through mirror transmission)
- 21 inches in front of NPRO (maybe junk b/c of ND filter? Maybe not?)
- Downstream of: HWP and PBS to get 120 mW total (about 8% power), and some focusing lenses (this was at PMC waist). I think this is the "junk light" you get from an NPRO. Maybe this is a lot of power to have in junk, maybe not?
|
Attachment 1: H1NPROoutput.PNG
|
|
Attachment 2: H1NPROat21in.PNG
|
|
Attachment 3: H1NPRO120mW.PNG
|
|
908
|
Fri Aug 6 00:44:06 2010 |
Frank | Lab Infrastructure | General | wifi bridges configured |
see elog entry here : http://nodus.ligo.caltech.edu:8080/PSL_Lab/257 |
907
|
Wed Aug 4 23:19:19 2010 |
Alastair | Laser | GYRO | Locking.... |
Quote: |
After we took that noise spectrum we started putting in the PD for the transmission readout. It can only take 55uW according to the spec sheet so we started by turning the laser power down a bit and then we were putting in a BS to reduce the power even more. We were just getting this set up when the cavity stopped locking properly. I was working on the theory that it was the lower laser power was causing issues, so I turned it back up - still won't lock. I checked the beam is on the reflection PDs (it is) and that they weren't saturating (they're not).
We're getting big dips in power on the reflection PD when I scan the cavity through resonance. We also get a big error signal out of the mixer when I scan the cavity, but it looks kind of one sided (there is a massive positive side and hardly any negative side). I can also see the error signals from the sidebands resonating, and they look more normal.
What we're seeing when we try to lock is that it only stays locked for a couple of seconds, and we get big fluctuations in the control signal to the piezo, before it falls out of lock.
More investigation will ensue.
PS. The mojo beads were on the laser - Jenna has removed them in case they were giving it bad vibes.
|
Still not got the cavity back to locking properly yet. It'll have to wait till tomorrow. I don't understand why the PDH signal has a really large positive side and small negative side when you scan the cavity - it needs more investigating.
I popped out to the 40m around 9pm to check the Rb clocks. They had come out of lock earlier (possibly due to being too hot inside the igloo). Rana has put a post up here about them http://nodus.ligo.caltech.edu:8080/40m/3361
They take about 12hrs to lock so the estimate of tomorrow afternoon seems reasonable. |
906
|
Wed Aug 4 20:34:48 2010 |
Alastair | Laser | GYRO | AOM Feedback Signal |
Quote: |
That's great! Only a factor of 1 million to go!!
The Tektronix generator is a horrible, horrible, monster and it should not be used as a VCO if at all possible. What's the loop gain for the CW and CCW loops? We need measurements. Also need to see the error point spectra of these loops out to high frequency. By looking at the error spectrum before and after a small increase in gain, one can infer the UGF.
Also, the 40m has one spare PDA255 which can be used to measure the beat frequency in transmission. Its in a plastic tub next to the RF stuff. Luckily Zach is working on a resonant RFPD and we can have something real very soon.
|
A million seems like a nice round number to start from.
We found one photodiode over here that we can use for the transmission readout. I was putting it in.... then the stuff in that other elog post happened. |
905
|
Wed Aug 4 20:28:03 2010 |
Alastair | Laser | GYRO | Locking.... |
After we took that noise spectrum we started putting in the PD for the transmission readout. It can only take 55uW according to the spec sheet so we started by turning the laser power down a bit and then we were putting in a BS to reduce the power even more. We were just getting this set up when the cavity stopped locking properly. I was working on the theory that it was the lower laser power was causing issues, so I turned it back up - still won't lock. I checked the beam is on the reflection PDs (it is) and that they weren't saturating (they're not).
We're getting big dips in power on the reflection PD when I scan the cavity through resonance. We also get a big error signal out of the mixer when I scan the cavity, but it looks kind of one sided (there is a massive positive side and hardly any negative side). I can also see the error signals from the sidebands resonating, and they look more normal.
What we're seeing when we try to lock is that it only stays locked for a couple of seconds, and we get big fluctuations in the control signal to the piezo, before it falls out of lock.
More investigation will ensue.
PS. The mojo beads were on the laser - Jenna has removed them in case they were giving it bad vibes. |
904
|
Wed Aug 4 20:26:30 2010 |
rana | Laser | GYRO | AOM Feedback Signal |
That's great! Only a factor of 1 million to go!!
The Tektronix generator is a horrible, horrible, monster and it should not be used as a VCO if at all possible. What's the loop gain for the CW and CCW loops? We need measurements. Also need to see the error point spectra of these loops out to high frequency. By looking at the error spectrum before and after a small increase in gain, one can infer the UGF.
Also, the 40m has one spare PDA255 which can be used to measure the beat frequency in transmission. Its in a plastic tub next to the RF stuff. Luckily Zach is working on a resonant RFPD and we can have something real very soon.
|
903
|
Wed Aug 4 18:52:18 2010 |
Jenna | Laser | GYRO | AOM Feedback Signal |
Here's a plot of the AOM feedback signal. The SR560 is DC coupled with a gain of 10, and the calibration for the data is
6.1e-4 V/count / 10 x 700 kHz/V * lambda*S/4A * 2pi (rad/s)/Hz = 3.618e-4
6.1e-4 is the calibration from the DAQ, 10 is the gain from the SR560, 700kHz/V is the deviation set on the Techtronix function generator, S is the perimeter of the cavity, and A is the area. One side of the cavity is .7889m. |
Attachment 1: AOM8-4.pdf
|
|
902
|
Wed Aug 4 17:24:11 2010 |
Alastair | Laser | GYRO | Photodiode for transmission |
I just ordered a 1811 photodiode from Newport for temporary transmission readout duty. I've ordered the one with the AC coupled output. It's in stock so should be with us soon. |
901
|
Wed Aug 4 17:21:56 2010 |
Alastair | Laser | Doubling | Mode Matching for Ovens |
Sorry about the camera. I was using it and didn't put it back in the drawer. It's in there now.
Quote: |
Quote: |
Quote: |
I redid the mode matching into the ovens to account for the dispersion of the material affecting the Rayleigh range (like this)
I used Zr = L_crystal / 2 / n
|
Scratched in favor of the following:
f=100mm at 0.47 m
f=75mm at 0.758 m
This works well with the current table layout.
taking Zin at PMC to be 371^2*pi/1.064 um
and Zoven to be 1.8*10000/2 um, about 36-37 inches from the PMC
|
I put these mode matching lenses in for the ovens...I realigned the Mach Zehnder with no focusing lenses for the PDs and was able to see fringes on the PD.. I can't find the lab camera and my iPhone died, so no picture.
Without trying very hard, I was able to get fringes with a min of 2.32V and a max of 4.88V on the scope, in green. When I get back from LLO I will be going straight to phase noise spectra.
|
|
900
|
Tue Aug 3 23:15:14 2010 |
Dmass | Laser | Doubling | Mode Matching for Ovens |
Quote: |
Quote: |
I redid the mode matching into the ovens to account for the dispersion of the material affecting the Rayleigh range (like this)
I used Zr = L_crystal / 2 / n
|
Scratched in favor of the following:
f=100mm at 0.47 m
f=75mm at 0.758 m
This works well with the current table layout.
taking Zin at PMC to be 371^2*pi/1.064 um
and Zoven to be 1.8*10000/2 um, about 36-37 inches from the PMC
|
I put these mode matching lenses in for the ovens...I realigned the Mach Zehnder with no focusing lenses for the PDs and was able to see fringes on the PD.. I can't find the lab camera and my iPhone died, so no picture.
Without trying very hard, I was able to get fringes with a min of 2.32V and a max of 4.88V on the scope, in green. When I get back from LLO I will be going straight to phase noise spectra. |
899
|
Tue Aug 3 21:43:40 2010 |
Dmass | Laser | Doubling | Mode Matching for Ovens |
Quote: |
I redid the mode matching into the ovens to account for the dispersion of the material affecting the Rayleigh range (like this)
I used Zr = L_crystal / 2 / n
|
Scratched in favor of the following:
f=100mm at 0.47 m
f=75mm at 0.758 m
This works well with the current table layout.
taking Zin at PMC to be 371^2*pi/1.064 um
and Zoven to be 1.8*10000/2 um, about 36-37 inches from the PMC |
898
|
Tue Aug 3 17:27:20 2010 |
Dmass | Laser | PSL | H1 NPRO Measured |
These rayleigh ranges correspond to waists of 151 and 146 um.
The spec for the Mephisto from the operating manual which was shipped with the 35W system gives a location of 9 cm inside the laser, not terribly different from 5 inches inside the laser. They did not seem to provide a number for typical waist size. Contacting Rick. |
897
|
Tue Aug 3 13:47:55 2010 |
Dmass | Laser | PSL | H1 NPRO Measured |
Quote: |
I redid the NPRO measurement and got this:
This might explain some of the rediculousitude of my inability to get mode matching to make sense...
|
This explains something. The chronological story (elog entries hyperlinked):
- I measured the waist of the NPRO (either I did it wrong, or it changed with temp/current)
- I calculated the mode matching for that waist, with 250mm and 200mm lenses
- I put them on the table in what I calculated to be the right spot
- I measured the waist of the input beam via the transmission through a steering mirror at the PMC waist
- It was grossly wrong
- I adjusted the mirrors by hand to give it the right waist (call these positions X1 and X2)
- I aligned into the PMC, and locked it, and noticed I only had about 25% transmission
- I measured the PMC waist
- I tried to adjust the lens positions by hand, hoping that my mode matching solution was sensitive to position / focal length errors.
- I moved the lenses by about a half inch in a few directions, realigning common/differential steering into the PMC each time
- I noticed no appreciable gain in any direction
- This is not surpising because I set the position of the lenses by forcing a 371 um waist at the PMC waist location
- I looked at the mode coming out of the laser, and on the PMC reflection
- I despaired that I might just not have much power in my TEM00
- I measured the beam profile after my first mode matching lens as a sanity check (250 mm focal length at 18 in) - plot at bottom of this elog
- I noticed that I couldn't get fits which were all that good - does this mean my M^2 is high?
- I redid the mode matching into the PMC with new lenses, aligned them into the PMC, locked it, and noticed a donut mode on REFL
- This is indicative of my mode matching being terrible
- I measured the waist with the wincamD at the PMC, it was WAY off (2mm instead of 370 um)
- I redid the beam waist measurement of the NPRO
- It looks drastically different
- I have changed both laser current slightly, and had feedback to the temperature
- I dont know if the above could cause that huge of a mode shift, or if it's just a somehow failed measurement with the wincamD. It's NOT a factor of 2.
- It had a funny mode shape (to be attached)
- I redid the mode matching with the new NPRO waist, and old lenses, and got the attached plot
- This is very similar to the first set of positions which I got moving the lenses by hand and using the wincamD, which leads me to believe the mode matching was roughly fine in the first place.
- I expect to find similar crappy (~25%) power transmission with this (old/new) mode matching solution and the "new" beam waist measurement, as I think I already did just this on the table
- If this is the case, I wonder if using a different laser is better...
- I was thinking of doing a cavity scan to see if there was a lot of "junk"
- If I do a cavity scan, will I see all the modal power? - that is to say, if I have a lot of higher order content, will it necessarily resonate in the cavity at some length
- Maybe what I'm asking is "do the resonant mode of the cavity form a complete basis for the HG laser polynomials?"
- Intuitively it seems like I could just get "unlucky" and have a lot of power in misc. mode which happen to not resonate strongly / at all in the cavity when my 00 is aligned perfectly - I'm half expecting someone to respond and let me know that this is not possible, and if I lock the cavity and align to 00, then scan, I have to see the higher order content I have in transmitted power.
|
Attachment 1: Npro2PMCmm.pdf
|
|
896
|
Tue Aug 3 01:31:15 2010 |
Dmass | Laser | PSL | H1 NPRO Measured |
I redid the NPRO measurement and got this:
This might explain some of the rediculousitude of my inability to get mode matching to make sense... |
Attachment 1: H1NPROWaist2.pdf
|
|
895
|
Mon Aug 2 22:23:36 2010 |
Dmass | Laser | PSL | H1 NPRO Measured |
I had only 25% coupling through the PMC, so was double checking my mode matching with the wincamD.
- Measurement of the H1NPRO beam profile
- Proposed mode matching solution (updated)
- I double checked my mode matching solution with a program Frank put on fb1 (called mm.exe, under /home/controls) from one of the Germans. The results were consistent.
- The mode I get at the PMC is WAY OFF from what it should be. ~750-800 um waist radius...
- I saw this inconsistency on the first iteration of mode matching as well, and just tuned the mirrors by hand using the wincamD to make it 370 um and put it in the right place.
- I had about 25% coupling max with the previous iteration.
I took some profiles with the WINCAMD of the transmission through one of the steering mirrors, and noticed that my fit didn't seem to....fit. This seems consistent with having a high M^2...Some notes on the measurement:
- WinCam did not seems saturated (no white area on the CCD 2d color scale image)
- Bleeding (blooming?) on CCD not on either measurement axis (also was not large)
- used X orientation of the U and V axes for getting waists
IF I DOUBLE CHECK MY MEASUREMENT OF THE H1NPRO WAIST SIZE / LOCATION, AND CONFIRM IT, WHAT DOES THIS ALL MEAN FOR ME?
The fear: There is just loads of higher order content in the beam, and only 25% of the power is in the TEM00...
From Frank - if it's the alignment of the pump diodes inside the NPRO, we can see this quickly by opening up the NPRO. I do not desire to do this before I finish my paper/experiment.
|
Attachment 1: H1MM2PMC.pdf
|
|
894
|
Mon Aug 2 21:02:19 2010 |
Dmass | Lab Infrastructure | General | PMC Waist |
I had never measured the output of the PMC. Here is a measurement of the waist at low power (~120 mW) with the H1NPRO. Nota Bene: The input mode is somewhat trash at this power (pics to be included later). I used the vertical axis of the wincamD to take this (should be tilt insensitive) |
Attachment 1: PMCwaist.pdf
|
|